Flashlight batteries is having electro-motive force of 1.5 v and internal resistance of one owns are connected to a resistor

of 5 ohms calculate the current in this resistor if the cell are connected to series or parallel


1
Expert's answer
2021-12-06T09:49:08-0500

The current:


"I=\frac V{R+r}=\frac{1.5}{1+5}=0.25\text{ A}."